Barnaby’s Fatburner Gyms Ltd operates a chain of exercise facilities throughout Victoria and New South Wales. The firm is considering offering share dividends to its

shareholders either via share bonus issue or share split. The firm currently has 8 million shares outstanding, which have a current market price of $12. If all else remains

constant, what will be the price of Barnaby’s shares after each of the following scenarios:

a. A 20% bonus share issue

b. A 32.5% bonus share issue

c. A 4 for 1 share split

d. What would be the total number of shares outstanding after (a), (b) and (c)?

2. Green Curry Kitchen operates two restaurants in Hobart and Launceston, Tasmania,

and has the following financial structure:

Account payables $100,000
Short-term debt $400,00
Current liabilities $ 500,00
Long-term debt $2,000,000
owner's equity $1,500,000
Total $4,000,000

The firm is considering an expansion that would involve raising an additional $2 million.

a. What are the firm’s debt ratio and debt-to-equity ratio in its present capital structure?

b. If the firm wants to have a debt ratio of 50%, how much equity does it need to raise in order to finance the expansion?
Answer #1

I am treating all these scenarios INDEPENDENT from each other. Also, the main concept behind a bonus share or a stock split is that the market cap remains the same, which means that
MARKET CAP before bonus or split = MARKET CAP after bonus or split

Market cap before = 12 X 8 million shares = $96 million

(a) A 20% bonus share issue
Existing shares = 8 million
Bonus at 20% = 20% X 8 million = 1.6 million share
New number of shares = 9.6 million shares
Total market cap = $96 million

Price per share = Market cap / number of shares = 96/9.6 = $10 per share

(b) A 32.5% bonus share issue
Existing shares = 8 million
Bonus at 32.5% = 32.5% X 8 million = 2.6 million shares
New number of shares = 10.6 million shares
Total market cap = $96 million

Price per share = Market cap / number of shares = 96/10.6 = $9.06 per share

(c) A 4:1 split
Existing shares = 8 million
4 : 1 split = 4 X 8 million = 32 million shares
Total market cap = $96 million

Price per share = Market cap / number of shares = 96/32 = $3 per share

(d) Total outstanding share
a. 9.6 million
b. 10.6 million
c. 32 million

2
a. Debt ratio = Total Debt / Total Assets
We know that Total Assets = Total Debt + Equity
So Debt ratio = Total Debt / (Total Debt + Equity)
= 2,500,000 / (2,500,000 + 1,500,000)
= 62.5 %

b. Debt equity ratio = (Long term debt + Short term debt) / Equity
= (2,000,000 + 40,000) / 1,500,000 = 1.36 : 1

If the firm wants debt ratio of 50% it needs its debt to be equal to its equity.
Debt = $2,500,000
Equity = $1,500,000

It must issue fresh equity of $1,000,000